OpenStudy (anonymous):

please help!! What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2?

OpenStudy (anonymous):

y=mx+b 4=2(3)+b 4=6+b 4-6=b -2=b

OpenStudy (anonymous):

thank you!